A ball is thrown vertically upward, which is the positive direction. A little later it returns to its point of release. The ball is in the air for a total time of 13.2 s. What is its initial velocity? Neglect air resistance.
Answer #1

since total time given in the question = 13.20 s

so 2 u/g =13.2

u =(13.2*g)/ 2

= (13.2 *9.8) /2

= 64.68 m/s